OpenStudy (anonymous):

Find d^2y/dx^2 in terms of x and y for y^2 = 4x

OpenStudy (anonymous):

\(y^2 = 4x\) Let us find out first derivative wrt to \(x\): \[2y \cdot \frac{dy}{dx} = 4 \cdot 1 \implies 2y \cdot \frac{dy}{dx} = 4 \implies y \frac{dy}{dx} = 2\] Getting this?

OpenStudy (anonymous):

So then the result is 2 ??

OpenStudy (anonymous):

We have not found second derivative.. :P

OpenStudy (anonymous):

\[y \frac{d^2y}{dx^2} + \frac{dy}{dx} \cdot (\frac{dy}{dx}) = 0 \implies y \frac{d^2y}{dx^2} = - (\frac{dy}{dx})^2 \color{blue}{ \implies \frac{d^2y}{dx^2} = -\frac{1}{y} \cdot (\frac{dy}{dx})^2}\]

OpenStudy (anonymous):

Mmm. how did you get 1/y ??

OpenStudy (anonymous):

Dividing both the sides by \(y\)..

OpenStudy (anonymous):

\(y\) on left hand side will cancel when we divide by \(y\) and right hand side, we will get : \[\frac{1}{y}\]

OpenStudy (anonymous):

ohhhh so is −1y⋅(dydx)^2 the final answer?

OpenStudy (anonymous):

-1/y * dy/dx the final result ???

OpenStudy (anonymous):

(dy/dx)^2 ??

OpenStudy (anonymous):

(-1/y) * (dy/dx)^2

OpenStudy (anonymous):

ok thank you for the help